A 1-kg ball moving at a velocity of 12 m/s collides head-on with a 2-kg ball moving at 24 m/s

in the opposite direction. Calculate the velocity of each ball after collision if (a) e = 2/3, (b) the

balls stick together, and (c) the collision is perfectly elastic.


1
Expert's answer
2021-03-14T19:18:51-0400

Answer

Given data

"u_1=24m\/s"

"u_1=-12m\/s"

"m_1=2kg\\m_2=1kg"

Initial momentum

"P_i=m_1u_1+m_2u_2=48-12\\=36kg-m/sec."

Velocitu after collosion

"v_1=\frac{P_i+em_2(u_2-u_1) }{m_1+m_2}"

"v_2=\frac{P_i+em_2(u_1-u_2) }{m_2+m_1}"

a)putting all value

e=2/3

"v_1=\frac{36+2/3\times1(-12-24) }{2+1}=4m/s"

"v_2=\frac{36+2/3\times1(24+12) }{2+1}=16m/s"

b) balls stick together

"P_i=(m_1+m_2) v"

"36=(2+1) v"

"v=12m\/sec"

c) perfectly elastic e=1

"v_1=\frac{36+2/3(-12-24) }{2+1}=0m/s"


"v_2=\frac{36+(24+12) }{2+1}=24m/s"